OpenStudy (anonymous):

whats this factorial 17!/15! ?

myininaya (myininaya):

\[\frac{17!}{15!}=\frac{17 \cdot 16 \cdot 15!}{15!}=17 \cdot 16 \]

OpenStudy (anonymous):

17x16=272.

OpenStudy (dumbcow):

\[\frac{17!}{15!} = \frac{17*16*15!}{15!} = 17*16\]

OpenStudy (anonymous):

\[17\times 16\]

OpenStudy (dumbcow):

im slow :)

myininaya (myininaya):

we all almost used different notation for multiplication

OpenStudy (anonymous):

guess this was a popular one

OpenStudy (anonymous):

17X16

OpenStudy (amistre64):

17.16